# Complete interpolating sequences for Fock type spaces

### 16 Mar 2026 v1

@@ 1. Introduction
A classical example is provided by the Paley–Wiener space of entire functions of exponential type. In this setting the Kadets–Ingham theorem describes the stability of exponential Riesz bases under small perturbations of the integer lattice (see [27]). More generally, in de Branges spaces complete interpolating sequences correspond exactly to Riesz bases of reproducing kernels [14]. These ideas have strongly influenced the modern theory of interpolation in spaces of analytic functions.
@@
In the setting of general sufficiently regular weights, one can describe conditions ensuring the existence of Riesz bases of reproducing kernels. For a weight $\varphi$, set $\psi(t)=2\varphi(e^{t}).$ Baranov, Belov and Borichev [5] proved that if $\psi^{\prime}(t)\to\infty$, $\psi^{\prime\prime}(t)>0$ is non-increasing, and
$|\psi^{\prime\prime\prime}(t)|=O\!\left((\psi^{\prime\prime}(t))^{5/3}\right),\qquad t\to\infty,$
then $\mathcal{F}^{2}_{\varphi}$ admits a Riesz basis of normalized reproducing kernels supported on the real line. The basis points are given by
@@ 1.1. Reference sequence and complete interpolation sequences for ${\mathcal{F}}^{p}_{\varphi}$
$\psi^{\prime}(t)\to\infty,\quad\psi^{\prime\prime}(t)>0\text{ is non-increasing, and }|\psi^{\prime\prime\prime}(t)|=O\bigl((\psi^{\prime\prime}(t))^{5/3}\bigr),\quad t\to\infty.$ (1.1)
@@
$\Sigma_{\varphi,p}=\{\sigma_{n}=e^{y_{n}}:\ \psi^{\prime}(y_{n})=pn+2,\ n\geq 0\}.$ (1.2)
@@
    $\Delta_{N}(\Gamma):=\limsup_{n\rightarrow\infty}\frac{1}{y_{n+N}-y_{n}}\left|\sum_{k=n+1}^{n+N}\delta_{k}\right|<\frac{1}{2}.$ (1.3)
@@ 1.2. Density results
${\mathcal{D}}_{\varphi}^{-}(\Gamma):=\liminf_{R\to\infty}\inf_{r\geq 0}\frac{n_{\Gamma}\bigl({\mathcal{C}}(e^{\eta(r)},e^{\eta(r+R)})\bigr)}{R},$ (1.4)
@@
${\mathcal{D}}_{\varphi}^{+}(\Gamma):=\limsup_{R\to\infty}\sup_{r\geq 0}\frac{n_{\Gamma}\bigl({\mathcal{C}}(e^{\eta(r)},e^{\eta(r+R)})\bigr)}{R}.$ (1.5)
@@
    If $\Gamma$ is a sampling set for ${\mathcal{F}}^{p}_{\varphi}$, then $\Gamma$ contains a subset $\Gamma^{\prime}$ such that ${\mathcal{D}}_{\varphi}^{-}(\Gamma^{\prime})\geq\frac{1}{p}$.
@@ 2. Preliminary Results
$\|f\|^{p}_{\varphi,p}\geq\int_{u-\varepsilon}^{u+\varepsilon}e^{\omega(t)-\psi(t)+2t}dt\gtrsim\ e^{-\psi(u)+u\psi^{\prime}(u)}\int_{u-\varepsilon}^{u+\varepsilon}e^{\omega(t)-\psi^{\prime}(u)t+2t}dt.$ (2.1)
@@
$\|z^{n}\|^{p}_{\varphi,p}\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{-1/2}e^{(pn+2)y_{n}-\psi(y_{n})},$ (2.2)
@@
$\left\|L_{e^{y_{n}}}\right\|\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2p}}e^{\frac{1}{p}\psi(y_{n})-\frac{2}{p}y_{n}},\quad n\geq 0.$ (2.3)
@@
$\|z^{n}\|^{p}_{\varphi,p}$ $=\int_{\mathbb{C}}|z|^{np}e^{-p\varphi(z)}dm(z)\asymp\int_{0}^{\infty}e^{(pn+2)t-\psi(t)}dt$
$\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{-1/2}e^{(pn+2)y_{n}-\psi(y_{n})}.$
@@
$\left\|L_{e^{y_{n}}}\right\|^{p}$ $\geq\frac{|e^{ny_{n}}|^{p}}{\|z^{n}\|^{p}_{\varphi,p}}\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2}}e^{pny_{n}}e^{-(np+2)y_{n}+\psi(y_{n})}$
$\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2}}e^{\psi(y_{n})-2y_{n}}.$
@@ 2.2. Canonical products estimates
$\sum_{k=1}^{m}y_{k}$ $=\sum_{k=1}^{m}(\psi^{\prime})^{-1}(pk+2)=\int_{0}^{m}(\psi^{\prime})^{-1}(pt+2)dt+\frac{y_{m}}{2}(1+o(1))$ (2.4)
@@
$=\frac{y_{m}}{p}\psi^{\prime}(y_{m})-\frac{\psi(y_{m})}{p}+O(1).$ (2.5)
@@
$\Gamma:=\left\{\gamma_{n}:=e^{y_{n}+\delta_{n}}e^{i\theta_{n}}\quad:\ n\geq 0\right\}.$
We associate with $\Gamma$ the canonical product
$G_{\Gamma}(z):=\prod_{\gamma\in\Gamma}\left(1-\frac{z}{\gamma}\right),\quad z\in\mathbb{C}.$ (2.6)
@@
$\Delta_{N}:=\limsup_{n\rightarrow\infty}\frac{1}{y_{n+N}-y_{n}}\left|\sum_{k=n+1}^{n+N}\delta_{k}\right|.$ (2.7)
@@
$\frac{1}{C}\frac{\mathrm{dist}(z,\Gamma)^{p}}{(1+|z|)^{2+\frac{p}{2}+p\Delta_{N}+\varepsilon}}\leq\left|G_{\Gamma}(z)\right|^{p}e^{-p\varphi(z)}\leq C\frac{\mathrm{dist}(z,\Gamma)^{p}}{(1+|z|)^{2+\frac{p}{2}-p\Delta_{N}-\varepsilon}},\quad z\in\mathbb{C},$ (2.8)
@@
$\frac{1}{C}\frac{1}{(1+|\gamma|)^{2+\frac{p}{2}+p\Delta_{N}+\varepsilon}}\leq\left|G^{\prime}_{\Gamma}(\gamma)\right|^{p}e^{-p\varphi(\gamma)}\leq C\frac{1}{(1+|\gamma|)^{2+\frac{p}{2}-p\Delta_{N}-\varepsilon}},\quad\gamma\in\Gamma.$ (2.9)
@@
$=pmt-p\sum_{n=0}^{m}y_{n}-p\sum_{n=0}^{m}\delta_{n}+p\log\mathrm{dist}(z,\Gamma)+O(1).$ (2.10)
@@
$=\left(\Delta_{N}+o(1)\right)t.$ (2.12)
@@
$\frac{\mathrm{dist}(z,\Gamma)^{p}}{(1+|z|)^{2+\frac{p}{2}+p\Delta_{N}+\varepsilon}}\lesssim\left|G_{\Gamma}(z)\right|^{p}e^{-p\varphi(z)}\lesssim\frac{\mathrm{dist}(z,\Gamma)^{p}}{(1+|z|)^{2+\frac{p}{2}-p\Delta_{N}-\varepsilon}},\quad z\in\mathbb{C}.$ (2.13)
@@
The following theorem extend the result obtained in [5, Theorem 1.2] and give similar result as in [22, Theorem 2.6]
@@
$f_{v}(z):=\sum_{n}v_{n}\|L_{\sigma_{n}}\|\frac{G_{\Sigma_{p}}(z)}{G^{\prime}_{\Sigma_{p}}(\sigma_{n})(z-\sigma_{n})},\quad z\in\mathbb{C}.$ (3.2)
@@
$|G_{\Sigma_{p}}(z)|^{p}\asymp e^{\ell(t)-2t}\mathrm{dist}(z,\Sigma_{p})^{p}.$ (3.3)
@@
$|G^{\prime}_{\Sigma_{p}}(\sigma_{n})|^{p}\asymp e^{\ell(y_{n})-2y_{n}},\quad n\geq 0.$ (3.4)
@@
$\left|\|L_{\sigma_{n}}\|\frac{G_{\Sigma_{p}}(z)}{G^{\prime}_{\Sigma_{p}}(\sigma_{n})(z-\sigma_{n})}\right|^{p}$ $\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2}}e^{\psi(y_{n})-\ell(y_{n})}e^{\ell(t)-2t}\frac{\mathrm{dist}(z,\Sigma_{p})^{p}}{|z-\sigma_{n}|^{p}}.$ (3.5)
@@
$|f_{v}(z)|e^{-\varphi(z)}\lesssim\sum_{n}|v_{n}|\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2p}}e^{\frac{1}{p}\psi(y_{n})-\frac{1}{p}\ell(y_{n})}e^{\frac{1}{p}\ell(t)-\frac{1}{p}\psi(t)-\frac{2}{p}t}\frac{\mathrm{dist}(z,\Sigma_{p})}{|z-\sigma_{n}|},\quad z\in\mathbb{C}.$ (3.6)
@@
$I(z)=\sum_{n}\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2p}}e^{\frac{1}{p}\psi(y_{n})-\frac{1}{p}\ell(y_{n})}\frac{\mathrm{dist}(z,\Sigma_{p})}{|z-\sigma_{n}|}\lesssim\left(\psi^{\prime\prime}(t)\right)^{\frac{1}{2p}}e^{\frac{1}{p}\psi(t)-\frac{1}{p}\ell(t)}.$ (3.7)
@@
$\lesssim\sum_{n}|v_{n}|^{p}\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2p}}e^{\frac{1}{p}\psi(y_{n})-\frac{1}{p}\ell(y_{n})}e^{\ell(t)-\psi(t)-2t}\frac{\mathrm{dist}(z,\Sigma_{p})}{|z-\sigma_{n}|}I(z)^{p-1}.$ (3.8)
@@
$=J_{n}^{(1)}+J_{n}^{(2)}\lesssim\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{-1}{2p}}e^{\frac{1}{p}\ell(y_{n})-\frac{1}{p}\psi(y_{n})}.$ (3.9)
@@
$f_{v}(z)=\sum_{n\geq-1}v_{n}e^{\varphi(\sigma_{n})}\frac{G(z)}{G^{\prime}(\sigma_{n})(z-\sigma_{n})},\quad z\in\mathbb{C},$ (3.10)
@@
$\|f\|_{\Sigma_{\varphi,p}}^{p}:=\sum_{n\geq 0}\ \left|\frac{f(\sigma_{n})}{\|L_{\sigma_{n}}\|}\right|^{p}$ (3.11)
@@ 4. Asymptotic evaluation estimates
$\asymp\sup_{\|(a_{n})\|_{p}\leq 1}\left|\sum_{n}a_{n}\frac{\|L_{\sigma_{n}}\|}{G_{\Sigma_{p}}^{\prime}(\sigma_{n})}\frac{G_{\Sigma_{p}}(z)}{z-\sigma_{n}}\right|.$ (4.1)
@@
$\|L_{z}\|$ $\asymp$ $\left(\sum_{n}\frac{\|L_{\sigma_{n}}\|^{q}}{|G_{\Sigma_{p}}^{\prime}(\sigma_{n})|^{q}}\frac{|G_{\Sigma_{p}}(z)|^{q}}{|z-\sigma_{n}|^{q}}\right)^{1/q}=S(z)^{1/q}.$ (4.2)
@@
$\langle u,v\rangle:=\sum_{n=0}^{\infty}u_{n}\overline{v_{n}},\quad u\in\ell^{p},\ v\in\ell^{\infty}.$ (4.3)
@@
$\|L_{z}\|\asymp\sup_{\|(a_{n})\|_{p}\leq 1}\left|\sum_{n}a_{n}\frac{\|L_{\sigma_{n}}\|}{G_{\Sigma_{p}}^{\prime}(\sigma_{n})}\frac{G_{\Sigma_{p}}(z)}{z-\sigma_{n}}\right|=\sup_{n}\frac{\|L_{\sigma_{n}}\|}{|G_{\Sigma_{p}}^{\prime}(\sigma_{n})|}\frac{|G_{\Sigma_{p}}(z)|}{|z-\sigma_{n}|}.$ (4.4)

### 23 Mar 2026 v2

@@ 1. Introduction
A classical example is provided by the Paley–Wiener space of entire functions of exponential type. In this setting the Kadets–Ingham theorem describes the stability of exponential Riesz bases under small perturbations of the integer lattice (see [16, 28]). More generally, in de Branges spaces complete interpolating sequences correspond exactly to Riesz bases of reproducing kernels [14]. These ideas have strongly influenced the modern theory of interpolation in spaces of analytic functions.
@@
In the setting of general sufficiently regular weights, one can describe conditions ensuring the existence of Riesz bases of reproducing kernels. For a weight $\varphi$, set $\psi(t)=2\varphi(e^{t}).$ Baranov, Belov and Borichev [5] proved that if $\psi$ satisfies suitable regularity conditions (see (1.1) below), then $\mathcal{F}^{2}_{\varphi}$ admits a Riesz basis of normalized reproducing kernels supported on the real line. The basis points are given by
@@ 1.1. Reference sequence and complete interpolation sequences for ${\mathcal{F}}^{p}_{\varphi}$
(1.1) $\psi^{\prime}(t)\to\infty,\quad\psi^{\prime\prime}(t)>0\text{ is non-increasing, and }|\psi^{\prime\prime\prime}(t)|=O\bigl((\psi^{\prime\prime}(t))^{5/3}\bigr),\quad t\to\infty.$
@@
(1.2) $\Sigma_{\varphi,p}=\{\sigma_{n}=e^{y_{n}}:\ \psi^{\prime}(y_{n})=pn+2,\ n\geq 0\}.$
@@
    (1.3) $\Delta_{N}(\Gamma):=\limsup_{n\rightarrow\infty}\frac{1}{y_{n+N}-y_{n}}\left|\sum_{k=n+1}^{n+N}\delta_{k}\right|<\frac{1}{2}.$
@@ 1.2. Density results
(1.4) ${\mathcal{D}}_{\varphi}^{-}(\Gamma):=\liminf_{R\to\infty}\inf_{r\geq 0}\frac{n_{\Gamma}\bigl({\mathcal{C}}(e^{\eta(r)},e^{\eta(r+R)})\bigr)}{R},$
@@
(1.5) ${\mathcal{D}}_{\varphi}^{+}(\Gamma):=\limsup_{R\to\infty}\sup_{r\geq 0}\frac{n_{\Gamma}\bigl({\mathcal{C}}(e^{\eta(r)},e^{\eta(r+R)})\bigr)}{R}.$
@@
    If $\Gamma$ is a sampling set for ${\mathcal{F}}^{p}_{\varphi}$, then ${\mathcal{D}}_{\varphi}^{-}(\Gamma)\geq\frac{1}{p}$.
@@ 2. Preliminary Results
(2.1) $\|f\|^{p}_{\varphi,p}\geq\int_{u-\varepsilon}^{u+\varepsilon}e^{\omega(t)-\psi(t)+2t}dt\gtrsim\ e^{-\psi(u)+u\psi^{\prime}(u)}\int_{u-\varepsilon}^{u+\varepsilon}e^{\omega(t)-\psi^{\prime}(u)t+2t}dt.$
@@
(2.2) $\|z^{n}\|^{p}_{\varphi,p}\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{-1/2}e^{(pn+2)y_{n}-\psi(y_{n})},$
@@
(2.3) $\left\|L_{e^{y_{n}}}\right\|\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2p}}e^{\frac{1}{p}\psi(y_{n})-\frac{2}{p}y_{n}},\quad n\geq 0.$
@@
$\|z^{n}\|^{p}_{\varphi,p}=\int_{\mathbb{C}}|z|^{np}e^{-p\varphi(z)}dm(z)\asymp\int_{0}^{\infty}e^{(pn+2)t-\psi(t)}dt\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{-1/2}e^{(pn+2)y_{n}-\psi(y_{n})}.$
@@
$\left\|L_{e^{y_{n}}}\right\|^{p}$ $\geq\frac{|e^{ny_{n}}|^{p}}{\|z^{n}\|^{p}_{\varphi,p}}\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2}}e^{pny_{n}}e^{-(np+2)y_{n}+\psi(y_{n})}\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2}}e^{\psi(y_{n})-2y_{n}}.$
@@ 2.2. Canonical products estimates
(2.4) $\sum_{k=1}^{m}y_{k}$ $=\sum_{k=1}^{m}(\psi^{\prime})^{-1}(pk+2)=\int_{0}^{m}(\psi^{\prime})^{-1}(pt+2)dt+\frac{y_{m}}{2}(1+o(1))$
@@
(2.5) $=\frac{y_{m}}{p}\psi^{\prime}(y_{m})-\frac{\psi(y_{m})}{p}+O(1).$
@@
$\Gamma:=\left\{\gamma_{n}:=e^{y_{n}+\delta_{n}}e^{i\theta_{n}}\quad:\ n\geq 0\right\}$
We associate to $\Gamma$ the canonical product
(2.6) $G_{\Gamma}(z):=\prod_{\gamma\in\Gamma}\left(1-\frac{z}{\gamma}\right),\quad z\in\mathbb{C}.$
@@
(2.7) $\Delta_{N}:=\limsup_{n\rightarrow\infty}\frac{1}{y_{n+N}-y_{n}}\left|\sum_{k=n+1}^{n+N}\delta_{k}\right|.$
@@
(2.8) $\frac{1}{C}\frac{\mathrm{dist}(z,\Gamma)^{p}}{(1+|z|)^{2+\frac{p}{2}+p\Delta_{N}+\varepsilon}}\leq\left|G_{\Gamma}(z)\right|^{p}e^{-p\varphi(z)}\leq C\frac{\mathrm{dist}(z,\Gamma)^{p}}{(1+|z|)^{2+\frac{p}{2}-p\Delta_{N}-\varepsilon}},\quad z\in\mathbb{C},$
@@
(2.9) $\frac{1}{C}\frac{1}{(1+|\gamma|)^{2+\frac{p}{2}+p\Delta_{N}+\varepsilon}}\leq\left|G^{\prime}_{\Gamma}(\gamma)\right|^{p}e^{-p\varphi(\gamma)}\leq C\frac{1}{(1+|\gamma|)^{2+\frac{p}{2}-p\Delta_{N}-\varepsilon}},\quad\gamma\in\Gamma.$
@@
(2.10) $=pmt-p\sum_{n=0}^{m}y_{n}-p\sum_{n=0}^{m}\delta_{n}+p\log\mathrm{dist}(z,\Gamma)+O(1).$
@@
(2.12) $=\left(\Delta_{N}+o(1)\right)t.$
@@
(2.13) $\frac{\mathrm{dist}(z,\Gamma)^{p}}{(1+|z|)^{2+\frac{p}{2}+p\Delta_{N}+\varepsilon}}\lesssim\left|G_{\Gamma}(z)\right|^{p}e^{-p\varphi(z)}\lesssim\frac{\mathrm{dist}(z,\Gamma)^{p}}{(1+|z|)^{2+\frac{p}{2}-p\Delta_{N}-\varepsilon}},\quad z\in\mathbb{C}.$
@@
The following theorem extend the result obtained in [5, Theorem 1.2] and give similar result as in [23, Theorem 2.6]
@@
(3.2) $f_{v}(z):=\sum_{n}v_{n}\|L_{\sigma_{n}}\|\frac{G_{\Sigma_{p}}(z)}{G^{\prime}_{\Sigma_{p}}(\sigma_{n})(z-\sigma_{n})},\quad z\in\mathbb{C}.$
@@
(3.3) $|G_{\Sigma_{p}}(z)|^{p}\asymp e^{\ell(t)-2t}\mathrm{dist}(z,\Sigma_{p})^{p}.$
@@
(3.4) $|G^{\prime}_{\Sigma_{p}}(\sigma_{n})|^{p}\asymp e^{\ell(y_{n})-2y_{n}},\quad n\geq 0.$
@@
(3.5) $\left|\|L_{\sigma_{n}}\|\frac{G_{\Sigma_{p}}(z)}{G^{\prime}_{\Sigma_{p}}(\sigma_{n})(z-\sigma_{n})}\right|^{p}$ $\asymp\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2}}e^{\psi(y_{n})-\ell(y_{n})}e^{\ell(t)-2t}\frac{\mathrm{dist}(z,\Sigma_{p})^{p}}{|z-\sigma_{n}|^{p}}.$
@@
(3.6) $|f_{v}(z)|e^{-\varphi(z)}\lesssim\sum_{n}|v_{n}|\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2p}}e^{\frac{1}{p}\psi(y_{n})-\frac{1}{p}\ell(y_{n})}e^{\frac{1}{p}\ell(t)-\frac{1}{p}\psi(t)-\frac{2}{p}t}\frac{\mathrm{dist}(z,\Sigma_{p})}{|z-\sigma_{n}|},\quad z\in\mathbb{C}.$
@@
(3.7) $I(z)=\sum_{n}\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2p}}e^{\frac{1}{p}\psi(y_{n})-\frac{1}{p}\ell(y_{n})}\frac{\mathrm{dist}(z,\Sigma_{p})}{|z-\sigma_{n}|}\lesssim\left(\psi^{\prime\prime}(t)\right)^{\frac{1}{2p}}e^{\frac{1}{p}\psi(t)-\frac{1}{p}\ell(t)}.$
@@
(3.8) $\lesssim\sum_{n}|v_{n}|^{p}\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{1}{2p}}e^{\frac{1}{p}\psi(y_{n})-\frac{1}{p}\ell(y_{n})}e^{\ell(t)-\psi(t)-2t}\frac{\mathrm{dist}(z,\Sigma_{p})}{|z-\sigma_{n}|}I(z)^{p-1}.$
@@
(3.9) $=J_{n}^{(1)}+J_{n}^{(2)}\lesssim\left(\psi^{\prime\prime}(y_{n})\right)^{\frac{-1}{2p}}e^{\frac{1}{p}\ell(y_{n})-\frac{1}{p}\psi(y_{n})}.$
@@
(3.10) $f_{v}(z)=\sum_{n\geq-1}v_{n}e^{\varphi(\sigma_{n})}\frac{G(z)}{G^{\prime}(\sigma_{n})(z-\sigma_{n})},\quad z\in\mathbb{C},$
@@
(3.11) $\|f\|_{\Sigma_{\varphi,p}}^{p}:=\sum_{n\geq 0}\ \left|\frac{f(\sigma_{n})}{\|L_{\sigma_{n}}\|}\right|^{p}$
@@ 4. Asymptotic evaluation estimates
In this section, we obtain asymptotic evaluation estimates for ${\mathcal{F}}^{p}_{\varphi}$, the linear maps $L_{z}\ :\ f\in{\mathcal{F}}^{p}_{\varphi}\longmapsto f(z)\in\mathbb{C}$. In particular, for $p=2$, we recover the known diagonal asymptotics of the reproducing kernel for weights of the form $\varphi(z)=(\log^{+}|z|)^{\alpha}$, $1<\alpha\leq 2$, obtained in [4, 17].
@@
(4.1) $\asymp\sup_{\|(a_{n})\|_{p}\leq 1}\left|\sum_{n}a_{n}\frac{\|L_{\sigma_{n}}\|}{G_{\Sigma_{p}}^{\prime}(\sigma_{n})}\frac{G_{\Sigma_{p}}(z)}{z-\sigma_{n}}\right|.$
@@
(4.2) $\|L_{z}\|$ $\asymp$ $\left(\sum_{n}\frac{\|L_{\sigma_{n}}\|^{q}}{|G_{\Sigma_{p}}^{\prime}(\sigma_{n})|^{q}}\frac{|G_{\Sigma_{p}}(z)|^{q}}{|z-\sigma_{n}|^{q}}\right)^{1/q}=S(z)^{1/q}.$
@@
(4.3) $\langle u,v\rangle:=\sum_{n=0}^{\infty}u_{n}\overline{v_{n}},\quad u\in\ell^{p},\ v\in\ell^{\infty}.$
@@
(4.4) $\|L_{z}\|\asymp\sup_{\|(a_{n})\|_{p}\leq 1}\left|\sum_{n}a_{n}\frac{\|L_{\sigma_{n}}\|}{G_{\Sigma_{p}}^{\prime}(\sigma_{n})}\frac{G_{\Sigma_{p}}(z)}{z-\sigma_{n}}\right|=\sup_{n}\frac{\|L_{\sigma_{n}}\|}{|G_{\Sigma_{p}}^{\prime}(\sigma_{n})|}\frac{|G_{\Sigma_{p}}(z)|}{|z-\sigma_{n}|}$